What is the common point between lines y = -x + 7 and y = 3x + 10?
1
Expert's answer
2012-08-07T11:25:00-0400
To find common point we should equate the expressions in the right parts:
& -x + 7 = 3x + 10
+x & +x

7=4x+10
-10 -10

-3=4x
whence x=-3/4
Corresponding value of y we can find by substituting obtained x=-3/4 into any of the equations:
y=-(-3/4)+7=3/4+7=31/4